# OmniUMI: Towards Physically Grounded Robot Learning via Human-Aligned Multimodal Interaction

### 12 Apr 2026 v1

@@ 4.4 Summary
Across all experiments, OmniUMI demonstrates reliable physically grounded sensing, improved human-aligned demonstration quality, and strong downstream learning performance. Accurate gravity compensation provides reliable external-wrench baselines, while the unified sensing and embodiment design enables high-fidelity multimodal data collection at scale. Compared with conventional teleoperation and ablated feedback settings, OmniUMI yields more stable interaction signals and more human-like demonstrations. Downstream evaluations further show that the proposed framework effectively supports grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control. Taken together, these results establish OmniUMI as a practical foundation for physically grounded robot learning via human-aligned multimodal interaction beyond vision-only representations.
@@ 5 Conclusion
We presented OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. Unlike prior UMI-like systems that primarily focus on visuomotor recording, OmniUMI moves toward interaction acquisition by jointly capturing tactile sensing, internal grasping force, external interaction wrench, and visual-motion observations within a single handheld platform, while organizing their acquisition around a human-aligned demonstration interface.
@@
The proposed framework is built around three tightly coupled components: a unified multimodal handheld interface for physically grounded data acquisition, a human-aligned interaction acquisition pipeline centered on dual-force feedback, and a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from robot-free multimodal human interaction to contact-rich robot execution.
@@
At the system and data-acquisition levels, OmniUMI supports practical multimodal integration, collection–deployment consistency, and physically grounded multimodal sensing within a shared embodiment. At the interaction-acquisition level, it improves human alignment during demonstration by enabling the operator to naturally perceive and regulate internal grasping force and external interaction wrench through dual-force feedback. At the learning and deployment levels, OmniUMI connects physically grounded and human-aligned multimodal data to downstream policy learning by extending diffusion policy with visual, tactile, and force-related observations, and translates the resulting policy outputs into controller-compatible execution through impedance-based deployment.
@@
Experiments show that OmniUMI improves sensing reliability, human-aligned demonstration quality, and downstream manipulation performance. In particular, the proposed framework supports three key interaction capabilities: grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control, validated respectively on force-sensitive pick-and-place, interactive surface erasing, and tactile-informed selective release. These results demonstrate that physically grounded multimodal interaction data can be reliably acquired, naturally regulated by human operators, and effectively leveraged for downstream contact-rich manipulation within a unified robot-free framework.
@@
At the same time, several limitations remain. Although the current results show improved signal-level alignment with human demonstration, a fuller causal account of how such alignment contributes to downstream policy performance is still lacking. In particular, while our experiments indicate that dual-force feedback improves the quality of collected interaction signals and benefits subsequent learning, the precise relationship between human-aligned interaction acquisition, data quality, and final task performance has not yet been isolated in a systematic manner. More broadly, the current study is evaluated on a limited set of contact-rich tasks and a single system instantiation, and a more comprehensive analysis across operators, embodiments, and task families remains necessary.
@@
These limitations point to several directions for future work. An important next step is to study human alignment more explicitly, for example through controlled ablations, user studies, and finer-grained analysis of how feedback design shapes demonstration behavior and downstream learning. It is also important to extend the framework to a broader range of embodiments and interaction settings, including more dexterous manipulation, longer-horizon tasks, and more diverse contact conditions. On the learning and control side, future work may further strengthen the coupling between multimodal policy learning and controller-compatible execution, including richer interaction representations and more principled impedance-aware deployment strategies.

### 22 Apr 2026 v2

@@ 4.4 Summary
Across all experiments, OmniUMI demonstrates reliable physically grounded sensing, improved human-aligned demonstration quality, and strong downstream learning performance. Accurate gravity compensation provides reliable external-wrench baselines, while the unified sensing and embodiment design enables high-fidelity multimodal data collection at scale. Compared with conventional teleoperation without bilateral force feedback, OmniUMI yields interaction signals that are consistently closer to direct human behavior across internal grasp force, external interaction wrench, and tactile deformation. Downstream evaluations further show that the proposed framework effectively supports grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control. Taken together, these results establish OmniUMI as a practical foundation for physically grounded robot learning via human-aligned multimodal interaction beyond vision-only representations.
@@ 5 Conclusion
We presented OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. Unlike prior UMI-like systems that primarily focus on visuomotor recording, OmniUMI moves toward interaction acquisition by jointly capturing tactile sensing, internal grasping force, external interaction wrench, and visual-motion observations within a single handheld platform, while organizing their acquisition around a human-aligned demonstration interface that supports natural perception and regulation of multimodal physical interaction.
@@
The proposed framework is built around three tightly coupled components: a unified multimodal handheld interface for physically grounded data acquisition, a human-aligned interaction acquisition pipeline grounded in bilateral gripper feedback and the handheld embodiment, and a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from robot-free multimodal human interaction to contact-rich robot execution.
@@
At the system and data-acquisition levels, OmniUMI supports practical multimodal integration, collection–deployment consistency, and physically grounded multimodal sensing within a shared embodiment. At the interaction-acquisition level, it improves human alignment during demonstration by enabling the operator to naturally perceive and regulate internal grasping force, external interaction wrench, and tactile interaction through bilateral gripper feedback and the handheld embodiment. At the learning and deployment levels, OmniUMI connects physically grounded and human-aligned multimodal data to downstream policy learning by extending diffusion policy with visual, tactile, and force-related observations, and translates the resulting policy outputs into controller-compatible execution through impedance-based deployment.
@@
Experiments show that OmniUMI improves sensing reliability, human-aligned demonstration quality, and downstream manipulation performance. In particular, the proposed framework supports three key interaction capabilities: grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control, validated respectively on force-sensitive pick-and-place, interactive surface erasing, and tactile-informed selective release. These results demonstrate that physically grounded multimodal interaction data can be reliably acquired, naturally perceived and regulated by human operators across internal grasp force, external contact force, and tactile interaction, and effectively leveraged for downstream contact-rich manipulation within a unified robot-free framework.
@@
At the same time, several limitations remain. Although the current results show improved signal-level alignment with human demonstration, a fuller causal account of how such alignment contributes to downstream policy performance is still lacking. In particular, while our experiments indicate that bilateral gripper feedback and the handheld embodiment improve the quality of collected interaction signals across internal grasp force, external contact force, and tactile interaction, and benefit subsequent learning, the precise relationship between human-aligned interaction acquisition, data quality, and final task performance has not yet been isolated in a systematic manner. More broadly, the current study is evaluated on a limited set of contact-rich tasks and a single system instantiation, and a more comprehensive analysis across operators, embodiments, and task families remains necessary.
@@
These limitations point to several directions for future work. An important next step is to study human alignment more explicitly, for example through controlled ablations, user studies, and finer-grained analysis of how bilateral feedback and handheld embodiment jointly shape internal-force, external-force, and tactile interaction during demonstration and downstream learning. It is also important to extend the framework to a broader range of embodiments and interaction settings, including more dexterous manipulation, longer-horizon tasks, and more diverse contact conditions. On the learning and control side, future work may further strengthen the coupling between multimodal policy learning and controller-compatible execution, including richer interaction representations and more principled impedance-aware deployment strategies.